### v2

# An extension of Steinberg’s Theorem to biquotient pairs of subgroups

An extension of Steinberg’s Theorem to biquotient pairs of subgroups
Marcus Zibrowius Research for this article was conducted within the framework of the DFG Research Training Group 2240: Algebro-Geometric Methods in Algebra, Arithmetic and Topology.
(October 22, 2024)
We study the derived tensor product of the representation rings of subgroups of a given compact Lie group $G$. That is, given two such subgroups $H_{1}$ and $H_{2}$, we study the tensor product of the associated representation rings $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ over the representation ring $\mathrm{R}G$, and prove a vanishing result for the associated higher Tor-groups. This result can be viewed as a natural extension of the Theorem of Steinberg that asserts that the representation rings of maximal rank subgroups of $G$ are free over $\mathrm{R}G$. It my also be viewed as an analogue of a result of Singhof on the cohomology of classifying spaces. We include an immediate application to the complex K-theory of biquotient manifolds.
Consider a compact connected Lie group $G$, with complex representation ring $\mathrm{R}G$. Given two closed connected subgroups $H_{1}$ and $H_{2}$, we can consider their representation rings $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ as $\mathrm{R}G$-modules. The purpose of this note is to prove:
Theorem 1.
Suppose the fundamental group of $G$ is torsion-free. If $H_{1}$ intersects every conjugate of $H_{2}$ trivially, then $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ vanishes for all $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})$.
This result is motivated by a theorem of Pittie and Steinberg, on which it builds and which it extends, and by a result of Singhof, of which it is an analogue. We briefly discuss each motivation in turn, and explain the implications for the complex K-theory of biquotient manifolds.
The first motivation is the following result which, following [Panin:twisted] and [Ananyevskiy:homogeneous], we simply refer as Steinberg’s Theorem. [^1] See Remark 17 for a historical discussion.
Theorem (Steinberg [Steinberg:Pittie]).
For any compact connected Lie group $G$ with torsion-free fundamental group, and for any closed connected subgroup $H$ of maximal rank (i. e. $\operatorname{rank}H=\operatorname{rank}G)$, the representation ring $\mathrm{R}H$ is free as an $\mathrm{R}G$-module.
Note that under the stated assumptions, a finitely generated $\mathrm{R}G$-module is free if and only if it is flat (see Lemma 16), so the theorem is equivalent to the assertion that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H,-)$ vanishes in all positive degrees $i$ for all connected subgroups $H$ of maximal rank. In particular – and this is historically the case of most interest – all higher Tor-groups $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H,\mathbb{Z})$ vanish for connected subgroups $H$ of maximal rank, which immediately implies that a certain spectral sequence of Hodgkin computing the complex K-theory of the homogeneous space $G/H$ collapses (see Section 4 for details). This case corresponds to the special case of Theorem 1 when $H_{1}$ is of maximal rank and $H_{2}$ is trivial, and it is in this sense that Theorem 1 is an extension.
The motivating result of Singhof can be stated as follows:
Theorem (Singhof [Singhof:DCM, Proposition (6.4)]).
Let $G$ be a compact connected Lie group (with arbitrary fundamental group), with closed connected subgroups $H_{1}$ and $H_{2}$. Consider the rational222 The theorem holds more generally for cohomology with coefficients in any integral domain for which the three cohomology rings $\mathrm{H}^{*}(\mathrm{B}G)$, $\mathrm{H}^{*}(\mathrm{B}H_{1})$ and $\mathrm{H}^{*}(\mathrm{B}H_{2})$ are polynomial algebras with generators in even dimensions. This is always true over $\mathbb{Q}$; see e. g. [McCleary, Theorem 6.38]. cohomology rings $\mathrm{H}^{*}(-):=\mathrm{H}^{*}(-;\mathbb{Q})$ of the classifying spaces $\mathrm{B}G$, $\mathrm{B}H_{1}$ and $\mathrm{B}H_{2}$. If $H_{1}$ intersects every conjugate of $H_{2}$ trivially, then the graded $\operatorname{Tor}$-groups $\operatorname{Tor}_{i}^{\mathrm{H}^{*}(\mathrm{B}G)}(\mathrm{H}^{*}(\mathrm{B}H_{1}),\mathrm{H}^{*}(BH_{2}))$ vanish for all $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})$.
We will refer to the condition on $H_{1}$ and $H_{2}$ appearing in Theorem 1 and in the result of Singhof as the strict biquotient condition. This condition arises rather naturally in nature. It is equivalent to the condition that $H_{1}$ acts freely from the left on the homogeneous space $G/H_{2}$ [Singhof:DCM, (1.1)], and hence implies that the double coset space, or biquotient, $H_{1}\backslash G/H_{2}$ is a smooth manifold. We thus refer to $H_{1}\backslash G/H_{2}$ as biquotient manifold.333 In [Singhof:DCM], the strict biquotient condition is referred to as ‘double coset condition’, and biquotient manifolds are called ‘double coset manifolds’. The term ‘biquotient’ used here appears to be vastly more popular.
The study of such biquotient manifolds has a rich history. We refer to [DeVito:thesis, Chapter 1] for a concise overview, in particular with respect to the quest for manifolds of positive curvature, and for an advertisement of biquotient manifolds as a fertile testing ground for conjectures relating the topology to the geometry of manifolds. For classification results, see in particular [Eschenburg:thesis, Ziller:lectures, KapovitchZiller, Totaro:biquotients, DeVito:6and7].
Singhof’s main application of the above vanishing result was a computation of the cohomology of biquotient manifolds, via an Eilenberg–Moore spectral sequence. Similarly, in Section 4, we will use Theorem 1 to compute the complex K-theory of biquotient manifolds, via the spectral sequence of Hodgkin already mentioned, generalizing the classical computation of the complex K-theory of homogeneous spaces:
Corollary 2.
Consider a compact connected Lie group $G$ with torsion-free fundamental group, closed connected subgroups $H_{1}$ and $H_{2}$ satisfying the strict biquotient condition, and the associated biquotient manifold $H_{1}\backslash G/H_{2}$. If $\operatorname{rank}H_{1}+\operatorname{rank}H_{2}\geq\operatorname{rank}G-1$, we have isomorphisms
$\mathrm{K}^{0}(H_{1}\backslash G/H_{2})$ $\cong\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\mathrm{R}H_{2}$
$\mathrm{K}^{1}(H_{1}\backslash G/H_{2})$ $\cong\operatorname{Tor}_{1}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$
In the maximal rank case, i. e. when $\operatorname{rank}H_{1}+\operatorname{rank}H_{2}=\operatorname{rank}G$, the groups in the second line vanish, i. e. $\mathrm{K}^{1}(G\backslash H_{1}/H_{2})=0$.
It follows, in particular, that under the stated assumptions the natural map from $\mathrm{R}(H_{1}\times H_{2})=\mathrm{R}H_{1}\otimes_{\mathbb{Z}}\mathrm{R}H_{2}$ to $\mathrm{K}^{0}(H_{1}\backslash G/H_{2})$ is surjective. This provides a partial answer to a question raised in [DeVitoGonzalez].
We will establish Theorem 1 in slightly greater generality than stated above.
First, the assumptions on $G$ can be relaxed slightly. Let’s say a Lie group $G$ is good if it is compact and connected and if $\mathrm{R}G$ is a polynomial ring tensored with a Laurent ring. All compact connected Lie groups with torsion-free fundamental group are good (Lemma 14), but not conversely. Steinberg established the theorem cited above under this more general assumption, see Theorem 13 below, and our Theorem 1 likewise holds for all good $G$.
Second, the strict biquotient condition can be relaxed as follows. Two subgroups $H_{1}$ and $H_{2}$ of $G$ satisfy the lax biquotient condition if every intersection of $H_{1}$ with a conjugate of $H_{2}$ is central in $G$. In this case, the action of $H_{1}$ on $G/H_{2}$ is still effectively free, and the double coset space $H_{1}\backslash G/H_{2}$ is still a smooth manifold. Theorem 1 also holds in this situation, but with an obvious correction term. Given any pair of subgroups $H_{1}$, $H_{2}$, we define the intersection rank $\mathrm{interrank}(H_{1},H_{2})$ as the maximum of the ranks of the intersections $H_{1}\cap gH_{2}g^{-1}$ as $g$ ranges over all elements of $G$. With this definition in place, the result reads as follows:
Theorem 3.
Let $G$ be a good Lie group, and let $H_{1}$ and $H_{2}$ be closed connected subgroups. If $H_{1}$ and $H_{2}$ satisfy the lax biquotient condition, then
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})=0$
for all $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})+\mathrm{interrank}(H_{1},H_{2})$.
We say a few words about the proof. There is little relation of our proof with Singhof’s proof, which makes crucial use of the grading in cohomology.
It turns out that the main difficulty is in establishing Theorem 3 in the maximal rank case, i. e. when $\operatorname{rank}G=\operatorname{rank}H_{1}+\operatorname{rank}H_{2}-\mathrm{interrank}(H_{1},H_{2})$. When $G$ is a torus, this case can be verified by a direct calculation using the spectral sequence of a double complex. The reader may easily reconstruct this calculation by stripping away all unnecessary complications from the proof of Proposition 32 below. It is also easy, for general $G$, to reduce to the case that $H_{1}$ and $H_{2}$ are tori; see Proposition 31. However, we have found no easy argument for passing from the case that the ambient group $G$ is a torus to the more general case. Instead, we try to salvage as much as possible from the calculation that is possible when $G$ is a torus. Two important ingredients are: (1) Segal’s theory of supports of prime ideals in representation rings [Segal:RG], which allows us to translate the biquotient condition into commutative algebra (Proposition 22), and (2) a “reduction to the diagonal”, which allows us to assume that $\mathrm{R}H_{2}$ is a quotient of $\mathrm{R}G$ by a regular sequence (see Proposition 25 and the proof of Proposition 33).
The passage from the maximal rank case to the general case is achieved by enlarging $H_{1}$ (see Proposition 29 and the proof of Proposition 34). However, in general this enlargment process destroys the biquotient condition (Remark 30). This forces us to prove Theorem 3 in somewhat greater generality than stated above, see Proposition 34. In short, we show that the Tor-groups always vanish at certain primes of $\mathrm{R}G$, and for subgroups satisfying the biquotient condition, these are the only primes that matter.
Acknowledgements
I am very grateful to Wilhelm Singhof for asking me whether Theorem 1 holds, and to David González-Álvaro and Panagiotis Konstantis for repeatedly motivating me to come up with an answer. Jeff Carlson and Oliver Goertsches shared their thoughts on various preliminary approaches to the problem. The special case of Theorem 1 in which $\operatorname{rank}H_{1}=1$ was established by Alessandra Wiechers in her Master’s thesis [Wiechers], already using Segal’s theory of supports. Leopold Zoller outlined a strategy to me by which the first isomorphism of Corollary 2 might alternatively be established in the case when $G$ is a product of special unitary and symplectic groups, and $H$ is of maximal rank, starting from the cohomological result of Singhof. Jakob Bergqvist shared some helpful insights regarding Lemma 11. Our use of the “reduction to the diagonal” is inspired by [Serre:LA, Chapter V.C, § 4].
1 Representation rings
We collect here assorted facts about complex representation rings of compact Lie groups that we will need later.
1.1 Abelian groups
For compact abelian Lie groups, the functor $\mathrm{R}$ that takes a compact abelian group to its complex representation ring can be factored into a composition of two functors as follows:
[…]
The first functor, which takes a compact abelian group $S$ to its character group ${\operatorname{Hom}(S,S^{1})}$, is a contravariant equivalence by Pontryagin duality. The second functor takes an abelian group $A$ to the associated group ring $\mathbb{Z}[A]$. In other words, $\mathrm{R}S\cong\mathbb{Z}[{\operatorname{Hom}(S,S^{1})}]$ for any compact Abelian group $S$. For example, when $T$ is a torus of rank $r$, i. e. when $T\cong(S^{1})^{r}$, we have $\operatorname{Hom}(T,S^{1})\cong\mathbb{Z}^{r}$. So $\mathrm{R}(T)$ is isomorphic to $\mathbb{Z}[x_{1}^{\pm 1},\dots,x_{r}^{\pm 1}]$, a Laurent ring on $r$ generators. We will need the following consequences of this explicit description.
Lemma 4.
Consider a torus $T$.
  1. (a)
    For an arbitrary closed subgroup $S\subseteq T$, the restriction $\mathrm{R}T\to\mathrm{R}S$ is surjective.
  2. (b)
    Any subtorus $S\subseteq T$ is a direct factor. For such a subtorus $S$, the kernel of the restriction $\mathrm{R}T\to\mathrm{R}S$ is generated by a regular sequence of length $\operatorname{rank}(T)-\operatorname{rank}(S)$.
  3. (c)
    There is an inclusion-reversing bijection between subgroups $S$ of $T$ and subgroups $K_{S}$ of $\mathbb{Z}^{r}$, where $r:=\operatorname{rank}T$, defined as follows: Identify ${\operatorname{Hom}(T,S^{1})}$ with $\mathbb{Z}^{r}$. Given a subgroup $S\subseteq T$, define the associated subgroup of $\mathbb{Z}^{r}$ as $K_{S}:=\ker(\operatorname{Hom}(T,S^{1})\to\operatorname{Hom}(S,S^{1}))$.
    Under this correspondence, $\operatorname{rank}(K_{S})=r-\operatorname{rank}(S)$, the subtori of $T$ correspond to the direct summands of $\mathbb{Z}^{r}$, and $K_{S_{1}\cap S_{2}}=K_{S_{1}}+K_{S_{2}}$.
1.2 Arbitrary compact groups
Definition 5.
We define the rank of a compact Lie group $G$ as the rank of any maximal torus contained in $G$.
Textbook sources such as [BtD:Lie, Adams:LieLectures]) restrict this definition to connected groups, but there is no harm in extending it to disconnected groups. Since any torus in $G$ is necessarily contained in the identity component of $G$, we are effectively defining the rank of $G$ as the rank of this connected component. In our main result, $G$ and the subgroups $H_{i}$ are assumed connected, but the intersections of $H_{1}$ with (conjugates of) $H_{2}$ may not be.
Lemma 6.
For any compact Lie group $G$:
  1. (a)
    $\mathrm{R}G$ is a free as an abelian group.
  2. (b)
    $\mathrm{R}G$ is finitely generated as a $\mathbb{Z}$-algebra, hence, in particular, a noetherian ring.
  3. (c)
    $\dim\mathrm{R}G=\operatorname{rank}G+1$
Proof.
The first observation follows directly from the definition of $\mathrm{R}G$ and is only mentioned here because we will implicitly use it when applying Theorem 20 later. For (b), see [Segal:RG, Corollary 3.3].
For (c), consider first a compact abelian group $S$. Pullback along the projection to the identity component $S\twoheadrightarrow S_{0}$ defines a finite ring extension $\mathrm{R}S_{0}\hookrightarrow\mathrm{R}S$. So $\dim(\mathrm{R}S)=\dim(\mathrm{R}S_{0})$. As $\mathrm{R}S_{0}=\mathbb{Z}[x_{1}^{\pm 1},\dots,x_{r}^{\pm 1}]$ with $r=\operatorname{rank}S_{0}=\operatorname{rank}S$, the claim follows in this case.
For any compact connected $G$, the restriction to a maximal torus $T$ defines an integral extension $\mathrm{R}G\to\mathrm{R}T$ (c. f. Lemma 8 below). So the result for $G$ follows from the result for $T$ which we have just established. For general compact $G$, we similarly have an integral extension $\mathrm{R}G\to\prod_{S}\mathrm{R}S$, where the product is over a finite number of compact abelian groups $S$ [Segal:RG, after Corollary (3.3)]. So $\dim\mathrm{R}G=\max_{S}\dim\mathrm{R}S$. More precisely, the groups $S$ range over all conjugacy classes of what Segal calls “Cartan subgroups” of $G$ (though note thas this terminology conflicts with modern usage). Not every such Cartan subgroup is of maximal rank, as the example in the final remark of [Segal:RG, § 1] shows. However, any maximal torus of the identity component of $G$ is a Cartan subgroup, so at least one of the groups $S$ has maximal rank, and the claim follows. ∎
Lemma 7.
For any subgroup $H$ of a compact Lie group $G$, $\mathrm{R}H$ is finitely generated as an $\mathrm{R}G$-module.
Proof.
This is [Segal:RG, Proposition 3.2]. ∎
1.3 Connected groups
We now restrict our attention to connected compact $G$. Recall that all maximal tori of $G$ are conjugate, and that any element of $G$ is contained in a maximal torus [Adams:LieLectures, Corollaries 4.22 & 4.23]. Given a maximal torus $T\subseteq G$, the Weyl group $\mathrm{W}:=\mathrm{W}(G,T)$ is defined as $\mathrm{W}:=\mathrm{N}_{G}T/T$. This is always a finite group. It acts on $\mathrm{R}T$ by conjugation, and we denote by $(\mathrm{R}T)^{\mathrm{W}}$ the subring of fixed points under this action.
Lemma 8.
Consider a compact connected Lie group $G$ with maximal torus $T$ and associated Weyl group $\mathrm{W}$. The restriction $i^{*}\colon\mathrm{R}G\to\mathrm{R}T$ is a finite extension which restricts to an isomorphism $\mathrm{R}G\cong(\mathrm{R}T)^{\mathrm{W}}$. In particular:
  1. (a)
    $\mathrm{R}G$ is an integral domain.
  2. (b)
    The set of prime ideals of $\mathrm{R}T$ lying over a fixed prime ideal $\mathfrak{p}\subset\mathrm{R}G$ is equal to a single $\mathrm{W}$-orbit of a prime in $\mathrm{R}T$.
  3. (c)
    A prime ideal $\mathfrak{p}_{T}\subset\mathrm{R}T$ lying over a given prime ideal $\mathfrak{p}\subset\mathrm{R}G$ is unique with this property if and only if it is $\mathrm{W}$-invariant. In this case, it contains every other prime ideal $\mathfrak{q}\subset\mathrm{R}T$ with $(i^{*})^{-1}\mathfrak{q}\subseteq\mathfrak{p}$, and $(\mathrm{R}T)_{\mathfrak{p}}$ is a local ring with maximal ideal $\mathfrak{p}_{T}\cdot(\mathrm{R}T)_{\mathfrak{p}}$.
Proof.
For the identification of $\mathrm{R}G$ with $(\mathrm{R}T)^{\mathrm{W}}$, see for example [Adams:LieLectures, Theorem 6.20]. For the finiteness of $i^{*}$, see Lemma 7 above. Alternatively, use the identification of $\mathrm{R}G$ with $(\mathrm{R}T)^{\mathrm{W}}$ and [GoertzWedhorn:AGI]*Proposition 12.27 (4).
Claim (a) is immediate from the fact that $\mathrm{R}T$ is an integral domain and the injectivity of $i^{*}$.
For claim (b), see for example [GoertzWedhorn:AGI]*Proposition 12.27 (2) and (3). The claim spells out the general fact that, for any commutative unital ring $A$, and any action of a finite group $\mathrm{W}$ through ring homomorphisms on $A$, the prime spectrum $\operatorname{Spec}(A^{\mathrm{W}})$ of the fixed ring can be naturally identified with $\operatorname{Spec}(A)/\mathrm{W}$, the set-theoretic quotient of the prime spectrum of $A$. (The same statement is true in the category of affine schemes, in fact more obviously so – see [GoertzWedhorn:AGI]*Proposition 12.27 (1). But it is the set-theoretic statement that we will need later.)
The first part of claim (c) follows immediately from (b). The second part follows from the first and the going-up property, which is shared by all finite (or, more generally, integral) extensions. Finally, note that the maximal primes of $(\mathrm{R}T)_{\mathfrak{p}}$ correspond to the primes of $\mathrm{R}T$ that lie over $\mathfrak{p}$. ∎
Example 9.
Let us write $\mathfrak{I}G$ for the kernel of the rank homomorphism $\operatorname{rank}\colon\mathrm{R}G\to\mathbb{Z}$. Clearly, $\mathfrak{I}T$ is a prime ideal of $\mathrm{R}T$ lying over $\mathfrak{I}G$, and it is $\mathrm{W}$-invariant. So Lemma 8 shows that $(\mathrm{R}T)_{\mathfrak{I}G}$ is a local ring with maximal ideal $\mathfrak{I}T\cdot(\mathrm{R}T)_{\mathfrak{I}G}$.
Similarly, for an arbitrary prime $p\in\mathbb{Z}$, we can consider the ideal $\mathfrak{I}_{p}G:=\operatorname{rank}^{-1}((p))$. Again, $\mathfrak{I}_{p}T$ is a $\mathrm{W}$-invariant prime ideal lying over $\mathfrak{I}_{p}G$, so $(\mathrm{R}T)_{\mathfrak{I}_{p}G}$ is a local ring with maximal ideal $\mathfrak{I}_{p}T\cdot(\mathrm{R}T)_{\mathfrak{I}_{p}G}$.
The inclusion of $\mathrm{R}G$ into $\mathrm{R}T$ moreover splits:
Theorem 10 (Holomorphic induction).
For any compact connected Lie group $G$ with maximal torus $i\colon T\hookrightarrow G$, we have a morphism of $RG$-modules $i_{*}\colon RT\to RG$ such that $i_{*}i^{*}=\operatorname{id}$.
Proof.
See [Atiyah:Bott]*Rmk 1 after Prop 4.9 or [EastwoodSalomon, § 3], in particular Lemma 3.4. ∎
Lemma 11.
The representation ring $\mathrm{R}G$ of any compact connected $G$ is biequidimensional, in the strong sense that all saturated chains of prime ideals have equal length, and hence satisfies the dimension formula: for any prime ideal $\mathfrak{p}\subset\mathrm{R}G$,
$\dim\mathrm{R}G=\dim(\mathrm{R}G/\mathfrak{p})+\dim(\mathrm{R}G)_{\mathfrak{p}}.$
Proof.
The dimensional formula holds in biquidimensional rings by [Heinrich, Proposition 4.1]. So we need only show that $\mathrm{R}G$ is biequidimensional. We refer to [Heinrich] for an in-depth discussion of this property. Suffice it to say that, in general, an equidimensional, coequidimensional and catenary ring is not necessarily biequidimensional. However, it is readily verified that the following three properties do imply biequidimensionality of a ring $R$:
  1. (i)
    $R$ has a unique minimal prime ideal,
  2. (ii)
    $R$ is catenary, i. e. all saturated chains of prime ideals between two given prime ideals in $R$ have equal length, and
  3. (iii)
    $R$ is equicodimensional, i. e. all its maximal ideals have equal height.
We now verify these two properties for $R=\mathrm{R}G$. Property (i) clearly holds as $\mathrm{R}G$ is an integral domain (Lemma 8 (a)). Property (ii) holds for $\mathrm{R}G$ as $\mathbb{Z}$ is universally catenary, as is any Dedekind domain, and as $\mathrm{R}G$ is finitely generated over $\mathbb{Z}$ (Lemma 6 (b)). Property (iii) holds for $R=\mathrm{R}G$ in view of the following general fact: any integral domain $R$ finitely generated over a Noetherian Jacobson universally catenary and equicodimensional domain $S$ is again equicodimensional [EGA4, Proposition 10.6.1 (i)]. This applies in particular to $S=\mathbb{Z}$, and more generally to any Dedekind domain $S$ with infinitely many maximal ideals. ∎
We include here another lemma that we will use in conjunction with Lemma 11.
Lemma 12.
Suppose $R\hookrightarrow S$ is a finite extension of noetherian rings, and that $\mathfrak{b}\subseteq S$ is some ideal. For any prime ideal $\mathfrak{p}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$, there exists a prime ideal $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$ such that $\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$ and such that
$\dim S_{\mathfrak{p}}-\dim(S/\mathfrak{b})_{\mathfrak{p}}=\dim R_{\mathfrak{p}}-\dim(R/\mathfrak{p}^{\prime})_{\mathfrak{p}}$
Proof.
For the ideal $\mathfrak{a}:=R\cap\mathfrak{b}$, we have a finite extension $R/\mathfrak{a}\hookrightarrow S/\mathfrak{b}$, and hence a finite extension $(R/\mathfrak{a})_{\mathfrak{p}}\hookrightarrow(S/\mathfrak{)}_{\mathfrak{p}}$ for any prime ideal $\mathfrak{p}\subset R$. In particular, $(\dim R/\mathfrak{a})_{\mathfrak{p}}=(\dim S/\mathfrak{b})_{\mathfrak{p}}$, and similarly $\dim S_{\mathfrak{p}}=\dim R_{\mathfrak{p}}$. Next, note that $\mathrm{Supp}_{R}(S/\mathfrak{b})=\mathrm{Supp}_{R}(R/\mathfrak{a})=\{\mathfrak{p}\in\operatorname{Spec}(R)\mid\mathfrak{a}\subseteq\mathfrak{p}\}.$ Indeed, the first equality again follows from the observation that $(R/\mathfrak{a})_{\mathfrak{p}}\hookrightarrow(S/\mathfrak{b})_{\mathfrak{p}}$ is an extension, and the second equality is elementary. So the given prime ideal $\mathfrak{p}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$ contains $\mathfrak{a}$, and $\dim(R/\mathfrak{a})_{\mathfrak{p}}=\dim(R/\mathfrak{p}^{\prime})_{\mathfrak{p}}$ for some prime ideal $\mathfrak{p}^{\prime}$ of $R$ minimal among the primes satisfying $\mathfrak{a}\subseteq\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$. In particular, $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$ and $\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$. ∎
1.4 Good groups and Steinberg’s Theorem
For any compact connected Lie group $G$, the commutator subgroup $G^{\prime}$ is semisimple [Sepanski, Theorem 5.22 (b)]. Moreover, $G$ can be identified with a quotient group $G\cong(G^{\prime}\times T)/\Gamma$, where $T$ is a torus and $\Gamma$ is a finite abelian subgroup of $G^{\prime}\times T$ intersecting $G^{\prime}$ trivially [Sepanski, Theorem 5.22 (a)]. Following Steinberg, we therefore refer to $G^{\prime}$ as the semisimple component of $G$. The following is the main result of [Steinberg:Pittie].
Theorem 13 ([Steinberg:Pittie, Theorem 1.2]).
For a compact connected Lie group $G$, the following conditions are equivalent:
  1. (a)
    The semisimple component $G^{\prime}$ of $G$ is a direct product of simple groups, each simply-connected or isomorphic to $\mathrm{SO}(2r+1)$ for some $r$.
  2. (b)
    The representation ring $\mathrm{R}G$ is a tensor product of a polynomial ring and a Laurent ring, i. e. isomorphic to the ring $\mathbb{Z}[x_{1},\dots,x_{m},y_{1}^{\pm 1},\dots,y_{n}^{\pm 1}]$ for certain $m$ and $n$.
  3. (c)
    For every connected subgroup $H$ of $G$ of maximal rank, $\mathrm{R}H$ is free as an $\mathrm{R}G$-module.
To avoid endlessly repeating these somewhat intricate conditions, we simply define a good Lie group to be a compact connected Lie group $G$ that satisfies the equivalent conditions of Theorem 13. The result stated as Steinberg’s Theorem is just the implication $(a\Rightarrow c)$ of this theorem combined with the following observation:
Lemma 14.
Any compact connected Lie group with torsion-free fundamental group is good.
Proof.
Suppose $G$ is compact connected with torsion-free fundamental group. Writing $G$ as $(G^{\prime}\times T)/\Gamma$ as above, we find that the fundamental group of $G^{\prime}$ is also torsion-free. As $G^{\prime}$ is semisimple, this implies that $G^{\prime}$ is simply-connected, and hence a direct product of simply-connected simple groups. So $G$ satisfies condition (a) of Theorem 13. ∎
Characterization (b) in Theorem 13 implies:
Lemma 15.
The representation ring $\mathrm{R}G$ of any good group $G$ is regular.
Proof.
For good $G$, $\mathrm{R}G$ is a localization of a polyonimal ring over the integers, which is clearly noetherian and of finite global dimension. ∎
Characterization (b) also implies that, for good $G$, the notions of “projective dimension” and “weak/flat dimenension” are equivalent for finitely generated $\mathrm{R}G$-modules. Indeed:
Lemma 16.
For good $G$, a finitely generated $\mathrm{R}G$-module is free if and only if is projective if and only it is flat.
Proof.
[…]
Remark 17 (History).
The implication $(a\Rightarrow b)$ of Theorem 13 is fairly elementary, and the implication $(b\Rightarrow c)$, which is the main ingredient in the result stated as Steinberg’s Theorem in the introduction, was in fact proved slightly earlier by Pittie [Pittie:Homogeneous], modulo the above generalization of the Quillen–Suslin theorem. Thus, from today’s vantage point, the main contribution of Steinberg is a proof of the opposite implication: a precise characterization of those compact Lie groups that have the desired property (c). Moreover, Steinberg’s proof includes an explicit basis of $\mathrm{R}H$ as an $\mathrm{R}G$-module.
We can easily sketch Pittie’s short argument for the implication $(b\Rightarrow c)$ based on the results and observations already included above: For a maximal torus $T$ in a good group $G$, both $\mathrm{R}G$ and $\mathrm{R}T$ are regular noetherian rings (Lemma 15). As $\mathrm{R}T$ is in any case finitely generated as an $\mathrm{R}G$-module (Lemma 7), it thus follows from [Serre:LA, Chapter III.D, § 5, Theorem 13 (b)] that $\mathrm{R}T$ is a projective $\mathrm{R}G$-module. For a general subgroup of maximal rank $H$, we may assume that $H$ contains $T$, and use holomorphic induction (Theorem 10) to deduce that $\mathrm{R}H$ is a direct summand of $\mathrm{R}T$. In particular, $\mathrm{R}H$ is also projective as an $\mathrm{R}G$-module. Finally, as we have just seen in Lemma 16, this is today known to imply that $\mathrm{R}H$ is free.
While Theorem 13 only deals with subgroups of maximal rank, the following consequence for subgroups of arbitrary ranks was proved in [SCSQ, Corollary 3.10]:
Corollary 18.
Let $G$ be a good Lie group, and let $H\subset G$ be an arbitrary closed connected subgroup. Then the projective dimension of $\mathrm{R}H$ as an $\mathrm{R}G$-module is at most $\operatorname{rank}G-\operatorname{rank}H$.
We do not reproduce the short proof of Corollary 18 here, but we note that besides holomorphic induction (Theorem 10) the main ingredient is a change-of-rings theorem for Tor. We record here two such theorems that we will have occassion to use later.
Theorem 19 (Change of rings for Tor I).
Fix a morphism of commutative rings $\Lambda\to\Gamma$, a $\Lambda$-module $A$ and a $\Gamma$-module $C$:
${A}$${C}$${\Lambda}$${\Gamma}$
If $\operatorname{Tor}^{\Lambda}_{*}(A,\Gamma)$ vanishes in all positive degrees, then we have isomorphisms of $\Lambda$-modules $\operatorname{Tor}^{\Lambda}_{i}(A,C)\cong\operatorname{Tor}^{\Gamma}_{i}(A\otimes_{\Lambda}\Gamma,C)$ in all degrees $i$.
Proof.
See [CartanEilenberg, Prop. VI.4.1.1].∎
Theorem 20 (Change of rings for Tor II).
Consider commutative $\mathbb{Z}$-algebras $\Gamma$, $\Lambda$ and $\Sigma$ which are free as abelian groups. Suppose $A$ is a $\Gamma\otimes\Lambda$-module, $B$ is a $\Lambda\otimes\Sigma$-module and $C$ is a $\Sigma\otimes\Gamma$-module, as summarized in the following diagram:
${A}$${B}$${C}$${\Lambda}$${\Sigma}$${\Gamma}$
Suppose $\operatorname{Tor}_{*}^{\Lambda}(A,B)$ and $\operatorname{Tor}_{*}^{\Sigma}(B,C)$ vanish in all positive degrees. Then we have natural isomorphisms of $\Sigma\otimes\Gamma$-modules
$\operatorname{Tor}_{i}^{\Gamma\otimes\Sigma}(A\otimes_{\Lambda}B,C)\cong\operatorname{Tor}_{i}^{\Lambda\otimes\Gamma}(A,B\otimes_{\Sigma}C)$
in all degrees $i$.
Proof.
This is a special case of [CartanEilenberg, Chapter IX, Theorem 2.8]. Cartan and Eilenberg work with non-commutative $K$-projective algebras $\Lambda$, $\Gamma$, $\Sigma$ over a commutative ring $K$. We have specialized the statement to the case of $\mathbb{Z}$-free commutative $\mathbb{Z}$-algebras. Note that the isomorphism $r$ of [CartanEilenberg, Proposition 2.1], on which the theorem relies, is an isomorphism of $\Sigma\otimes\Gamma$-algebras in this commutative situation. ∎
When $\Lambda$ and $\Gamma$ are free as abelian groups, Theorem 19 can be seen as a special case of Theorem 20.
1.5 Supports
In [Segal:RG], Segal analyses the prime spectrum $\operatorname{Spec}(\mathrm{R}G)$ of an arbitrary compact Lie group $G$. In particular, Segal shows that, for each prime ideal $\mathfrak{p}\subset\mathrm{R}G$, there is a subgroup $S_{\mathfrak{p}}$, unique up to conjugation, which is minimal with respect to the property that $\mathfrak{p}$ is the preimage of a prime ideal of $\mathrm{R}(S_{\mathfrak{p}})$ under the restriction $\mathrm{R}G\to\mathrm{R}S_{\mathfrak{p}}$. This subgroup $S_{\mathfrak{p}}$, or rather any subgroup in this conjugacy class, is referred to as the support of $\mathfrak{p}$. The following proposition explains, in particular, how Segal’s “supporting subgroups” and the usual notion of “prime ideals supporting a module” are related.
Proposition 21 (Segal).
Let $G$ be a c ompact Lie group, $H$ a closed subgroup of $G$ and $\mathfrak{p}\in\operatorname{Spec}(\mathrm{R}G)$. The following statements are equivalent:
  1. (a)
    $\mathfrak{p}\in\operatorname{im}(\operatorname{Spec}(\mathrm{R}H)\to\operatorname{Spec}(\mathrm{R}G))$
  2. (b)
    $\mathfrak{p}$ contains $\ker(\mathrm{R}G\to\mathrm{R}H)$
  3. (c)
    $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H$
  4. (d)
    The support $S_{\mathfrak{p}}$ is conjugate to a subgroup of $H$.
Proof.
See [Segal:RG, Proposition 3.7 (iv)] (or [Wiechers, Proposition 110]). ∎
Proposition 22.
Let $G$ be a compact connected Lie group with maximal torus $T$.
  1. (a)
    For subgroups $H^{\prime}\subseteq H\subseteq G$, we have $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H^{\prime}\subseteq\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H.$
  2. (b)
    For any two subgroups $H_{1}$ and $H_{2}$,
    $\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\mathrm{R}H_{2})$ $=\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$
    $=\bigcup_{g\in G}\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(H_{1}\cap gH_{2}g^{-1}).$
  3. (c)
    Given a subgroup $H\subseteq G$ and a prime ideal $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H$,
    $\dim(\mathrm{R}G/\mathfrak{p})\leq\operatorname{rank}H+1.$
  4. (d)
    Let $\mathrm{Z}G$ denote the centre of $G$. If $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$, there is exactly one prime ideal of $\mathrm{R}T$ lying over $\mathfrak{p}$.
Proof.
Claim (a) is clear from the equivalence $(c\Leftrightarrow d)$ of Proposition 21.
The first equality in Claim (b) holds in general for finitely generated modules over a commutative ring. See [AtiyahMacDonald, Chapter 3, Exercise 19 (iv)] for the general fact, and Lemma 7 for finite generation in the case at hand. The second equality in Claim (b) again follows from the equivalence $(c\Leftrightarrow d)$ of Proposition 21.
For claim (c), recall from [Segal:RG] that the support $S_{\mathfrak{p}}$ is “cyclic”, i. e. containes an element $s$ whose powers are dense in $S_{\mathfrak{p}}$. This element $s$ is contained in some maximal torus of $G$, so we may as well assume $s\in T$ and hence $S_{\mathfrak{p}}\subseteq T$. The ideal $\mathfrak{p}\subset\mathrm{R}G$ is thus the inverse image of an ideal $\mathfrak{p}_{0}\subset\mathrm{R}(S_{\mathfrak{p}})$ under a restriction map that factors as
$\mathrm{R}G\hookrightarrow\mathrm{R}T\twoheadrightarrow\mathrm{R}(S_{\mathfrak{p}}).$
Recall that any inclusion of a subgroup into a torus induces a surjection on representation rings (Lemma 4 (a)). Writing $\mathfrak{p}_{T}\subset\mathrm{R}T$ for the intermediate preimage of $\mathfrak{p}_{0}$, we obtain induced morphisms of quotient rings as follows:
$\mathrm{R}G/\mathfrak{p}\hookrightarrow\mathrm{R}T/\mathfrak{p}_{T}\xrightarrow{\cong}\mathrm{R}(S_{\mathfrak{p}})/\mathfrak{p}_{0}$
Here, the first map is a finite extension (see Lemma 8), so $\dim\mathrm{R}G/\mathfrak{p}=\dim\mathrm{R}T/\mathfrak{p}_{T}$, and we deduce $\dim(\mathrm{R}G/\mathfrak{p})\leq\dim(\mathrm{R}S_{\mathfrak{p}})$. Moreover, by Lemma 6 (c), $\dim(\mathrm{R}S_{\mathfrak{p}})=\operatorname{rank}(S_{\mathfrak{p}})+1$, and as $S_{\mathfrak{p}}$ is conjugate to a subgroup of $H$ we have $\operatorname{rank}(S_{\mathfrak{p}})\leq\operatorname{rank}H$.
For claim (d), recall that the centre of $G$ is contained in any maximal torus. So the assumption implies that $\mathfrak{p}\subseteq\mathrm{R}G$ can be obtained as the inverse image of some ideal in $\mathrm{R}(\mathrm{Z}G)$ under a restriction map that factors as
$\mathrm{R}G\hookrightarrow\mathrm{R}T\twoheadrightarrow\mathrm{R}(\mathrm{Z}G).$
Let us again write $\mathfrak{p}_{T}$ for the intermediate preimage on $\mathrm{R}T$, which by construction lies over $\mathfrak{p}$. The restriction $RT\twoheadrightarrow\mathrm{R}(\mathrm{Z}G)$ clearly commutes with the action of the Weyl group on $\mathrm{R}T$, so the Weyl group acts trivially on $\mathfrak{p}\in\operatorname{Spec}(\mathrm{R}T)$. It therefore follows from Lemma 8 (b) that $\mathfrak{p}_{T}$ is the only ideal over $\mathfrak{p}$. ∎
Example 23.
Recall the defition of the prime ideals $\mathfrak{I}G$ and $\mathfrak{I}_{p}G$ from Example 9. If $(G,H_{1},H_{2})$ satisfy the strict biquotient condition, then the intersection $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$ consists exclusively of the primes $\mathfrak{I}G$ and $\mathfrak{I}_{p}G$ for the various primes $p\in\mathbb{Z}$. Indeed, proposition (b) identifies this intersection with $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(1)$, and by the equivalence $(b\Leftrightarrow c)$ of Proposition 21 this latter support consists precisely of the primes containing $\mathfrak{I}G$.
2 Biquotient pairs
In this section, we gather preliminary results on biquotients that facilitate certain reduction steps in the proof of Theorem 3. We begin by recalling the key definitions from the introduction.
Definition 24.
A triple $(G,H_{1},H_{2})$ consisting of a compact Lie group $G$ and two closed subgroups $H_{1}$ and $H_{2}$ of $G$ satisfies the strict biquotient condition if $H_{1}\cap gH_{2}g^{-1}=\{1\}$ for every $g\in G$. It satisfies the lax biquotient condition if $H_{1}\cap gH_{2}g^{-1}$ is central in $G$ for every $g\in G$. The intersection rank of $H_{1}$ and $H_{2}$ in $G$ is defined as
$\mathrm{interrank}(H_{1},H_{2}):=\max_{g\in G}(\operatorname{rank}{H_{1}\cap gH_{2}g^{-1}}).$
A biquotient manifold is a double coset space $H_{1}\backslash G/H_{2}$ associated with a triple satisfying the (strict or) lax biquotient condition, with its induced smooth structure.
2.1 Reduction to diagonal
A biquotient manifold $H_{1}\backslash G/H_{2}$ can equivalently be viewed as a quotient
$(H_{1}\times H_{2})\backslash\!\backslash G:=(H_{1}\times H_{2})\backslash(G\times G)/\Delta G,$
where $\Delta\colon G\hookrightarrow G\times G$ denotes the diagonal subgroup. Explicitly, the isomorphism is given by $g\mapsto(g,1)$, with inverse $g_{1}g_{2}^{-1}\mapsfrom(g_{1},g_{2})$.
Proposition 25.
Let $G$ be a compact Lie group with closed subgroups $H_{1}$ and $H_{2}$. The triple $(G,H_{1},H_{2})$ satisfies the lax/strict biquotient condition if and only if the triple $(G\times G,H_{1}\times H_{2},\Delta G)$ does. Irrespective of this condition, $\mathrm{interrank}(H_{1},H_{2})=\mathrm{interrank}(H_{1}\times H_{2},\Delta G)$, and we have canonical isomorphisms of abelian groups
$\operatorname{Tor}_{i}^{\mathrm{R}(G\times G)}(\mathrm{R}(H_{1}\times H_{2}),\mathrm{R}G)\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2}).$
in all degrees $i$. Likewise, for any prime ideal $\mathfrak{p}\subset\mathrm{R}G$, we have isomorphisms of abelian groups
$\operatorname{Tor}_{i}^{\mathrm{R}(G\times G)}(\mathrm{R}(H_{1}\times H_{2}),\mathrm{R}G)_{(\Delta^{*})^{-1}(\mathfrak{p})}\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}},$
in all degrees $i$. Finally,
if $\mathfrak{p}$ $\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}(\mathrm{Z}G)),$
then $(\Delta^{*})^{-1}(\mathfrak{p})$ $\in\mathrm{Supp}_{\mathrm{R}(G\times G)}\mathrm{R}(\mathrm{Z}(G\times G)).$
Proof.
For the first assertions, note that $(H_{1}\times H_{2})\cap(g_{1},g_{2})\Delta G(g_{1},g_{2})^{-1}$ is isomorphic to $H_{1}\cap g_{1}g_{2}^{-1}H_{2}g_{2}g_{1}^{-1}$ under projection to the first coordinate, and that $\mathrm{Z}(G\times G)=\mathrm{Z}G\times\mathrm{Z}G$.
The isomorphism of Tor-modules can be obtained from Theorem 20 by taking $\Gamma$, $\Lambda$, $\Sigma$, $A$, $B$ and $C$ as indicated in the following diagram:
[…]
The Tor-modules assumed to vanish in positive degrees in Theorem 20 are $\operatorname{Tor}_{*}^{\mathbb{Z}}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ and $\operatorname{Tor}_{*}^{\mathrm{R}G}(\mathrm{R}H_{2},\mathrm{R}G)$. In both cases, this vanishing is clear because at least one of the modules is free over the considered base ring. So we obtain isomorphisms of $\mathrm{R}G\otimes\mathrm{R}G$-modules
$\operatorname{Tor}_{i}^{\mathrm{R}G\otimes\mathrm{R}G}(\mathrm{R}H_{1}\otimes\mathrm{R}H_{2},\mathrm{R}G)\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$
in all degrees $i$. To simplify notation, let us temporarily write $G^{2}$ for $G\times G$. Recall that we have a natural isomorphisms $\mathrm{R}(H_{1}\times H_{2})\cong\mathrm{R}H_{1}\otimes\mathrm{R}H_{2}$ and, in particular, $\mathrm{R}(G^{2})\cong\mathrm{R}G\otimes\mathrm{R}G$. We can therefore rewrite the above isomorphisms in the form asserted in the proposition. Under the identification $\mathrm{R}(G^{2})\cong\mathrm{R}G\otimes\mathrm{R}G$, the restriction along $\Delta$ is simply the ring multiplication $\Delta^{*}\colon\mathrm{R}(G^{2})\twoheadrightarrow\mathrm{R}G$. The $\mathrm{R}(G^{2})$-module structure on the right side of the isomorphism is induced from the usual $\mathrm{R}G$-module structure via this epimomorphism $\Delta^{*}$. The second, localized claim follows immediately.
For the final claim, recall again that $\mathrm{Z}(G^{2})=\mathrm{Z}G\times\mathrm{Z}G$. The epimorphism $\Delta^{*}\colon\mathrm{R}(G^{2})\twoheadrightarrow\mathrm{R}G$ restricts to an epimorphism $\mathrm{R}(\mathrm{Z}(G^{2}))\to\mathrm{R}(\mathrm{Z}G)$, which we can view as an epimorphism of $\mathrm{R}(G^{2})$-modules. Localizing at $(\Delta^{*})^{-1}(\mathfrak{p})$ for some prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ yields an epimorphism $\mathrm{R}(\mathrm{Z}(G^{2}))_{(\Delta^{*})^{-1}(\mathfrak{p})}\to(\mathrm{R}(\mathrm{Z}G))_{\mathfrak{p}}$. So if $\mathrm{R}(\mathrm{Z}G)$ is supported at $\mathfrak{p}$, then $\mathrm{R}(\mathrm{Z}(G^{2}))$ must be supported at $(\Delta^{*})^{-1}(\mathfrak{p})$. ∎
2.2 The biquotient condition can be checked on finitely many intersections
By definition, the biquotient condition is a condition on infinitely many intersections $H_{1}\cap gH_{2}g^{-1}$. We show here that only finitely many intersections need to be taken into account, and that, similarly, the intersection rank of $H_{1}$ and $H_{2}$ can be computed from only finitely many intersections. In short, we can replace $H_{1}$ and $H_{2}$ by tori (Proposition 26), and we need only consider conjugates by elements of the Weyl group of $G$ (Proposition 28).
Proposition 26.
Consider a compact Lie group $G$. For any pair of closed connected subgroups $H_{1}$ and $H_{2}$, with maximal tori $T_{1}$ and $T_{2}$, respectively, $\mathrm{interrank}(H_{1},H_{2})=\mathrm{interrank}(T_{1},T_{2})$. Moreover, the triple $(G,H_{1},H_{2})$ satisfies the lax/strict biquotient condition if and only if the triple $(G,T_{1},T_{2})$ does.
Proof.
By definition, $\mathrm{interrank}(H_{1},H_{2})$ is the maximum over the ranks of all tori $S$ contained in the intersections $H_{1}\cap gH_{2}g^{-1}$, where $g$ ranges over the elements of $G$. Any such torus $S$ is conjugate to a subgroup of $T_{1}$ and a subgroup of $T_{2}$, hence conjugate to a subgroup of $T_{1}\cap g^{\prime}T_{2}g^{\prime-1}$ for some $g^{\prime}\in G$. This shows that the interranks agree. If $H_{1}$ and $H_{2}$ satisfy the lax biquotient condition, then clearly so do $T_{1}$ and $T_{2}$. Conversely, suppose $T_{1}$ and $T_{2}$ satisfy the lax biquotient condition, and suppose that $h\in H_{1}\cap gH_{2}g^{-1}$ for some $g\in G$. As any element of $H_{1}$ is conjugate to an element of $T_{1}$, and any element of $H_{2}$ is conjugate to an element of $T_{2}$, we find that $h=h_{1}t_{1}h_{1}^{-1}=gh_{2}t_{2}h_{2}^{-1}g^{-1}$ for certain $h_{i}\in H_{i}$ and $t_{i}\in T_{i}$. It follows that $t_{1}\in T_{1}\cap g^{\prime}T_{2}g^{\prime-1}$ for $g^{\prime}:=h_{1}^{-1}gh_{2}$, hence $t_{1}$ is central by our assumption on $T_{1}$ and $T_{2}$. So $h=h_{1}t_{1}h_{1}^{-1}=t_{1}$ is central. The argument for the strict condition is analogous. ∎
Suppose now that $G$ is compact and connected, and choose a maximal torus $T$ in $G$. The associated Weyl group $\mathrm{W}:=\mathrm{N}_{G}T/T$ is a finite group. Given a subgroup $S\subseteq T$, we write $w(S)$ for the conjugate $hSh^{-1}$ given by an arbitrary representative $h\in\mathrm{N}_{G}T$ of $w\in\mathrm{W}$.
Lemma 27.
Consider a compact connected Lie group $G$ with a maximal torus $T$ and associated Weyl group $\mathrm{W}$. Suppose $T_{1}$ and $T_{2}$ are subtori of $T$. For each $g\in G$, there exists a $w\in\mathrm{W}$ such that $T_{1}\cap gT_{2}g^{-1}\subseteq T_{1}\cap w(T_{2})$.
Proof.
It suffices to show that, for every element $g\in G$, there exists an element $w\in\mathrm{W}$ such that $T\cap gT_{2}g^{-1}\subseteq w(T_{2})$. We show this by adapting arguments from the proof of [BtD:Lie]*IV, Lemma 2.5. Consider $S:=T\cap gT_{2}g^{-1}$. Conjugation by $g$ restricts to an isomorphism between centralizers: $c(g)\colon\mathrm{Z}_{G}(g^{-1}Sg)\to\mathrm{Z}_{G}S$. As $g^{-1}Sg\subseteq T$, and as $T$ is abelian, $T\subseteq\mathrm{Z}_{G}(g^{-1}Sg)$, so $c(g)(T)\subseteq\mathrm{Z}_{G}S$. Likewise, $T\subseteq\mathrm{Z}_{G}S$. So $c(g)(T)$ and $T$ are two maximal tori contained in (the identity component of) $\mathrm{Z}_{G}S$, and hence conjugate in (the identity component of) $\mathrm{Z}_{G}S$. So there exists some element $f\in\mathrm{Z}_{G}S$ such that $fgTg^{-1}f^{-1}=T$, i. e. such that $fg\in\mathrm{N}_{G}T$. Take $w$ to be the image of $fg$ in $\mathrm{W}$. The equality $S=fSf^{-1}$ now expands to $T\cap gT_{2}g^{-1}=f(T\cap gT_{2}g^{-1})f^{-1}$, and the group on the right is clearly contained in $fgT_{2}g^{-1}f^{-1}=w(T_{2})$. ∎
The following proposition is immediate from Lemma 27.
Proposition 28.
Consider a compact connected Lie group $G$ with a maximal torus $T$ and associated Weyl group $\mathrm{W}$. For any pair of subtori $T_{1}$ and $T_{2}$ of $T$, $\mathrm{interrank}(T_{1},T_{2})=\max_{w\in\mathrm{W}}(\operatorname{rank}{T_{1}\cap w(T_{2})})$. Moreover, the triple $(G,T_{1},T_{2})$ satisfies the lax/strict biquotient condition if and only if the intersection $T_{1}\cap w(T_{2})$ is trivial/central for each $w\in\mathrm{W}$. ∎
Note that we can always arrange for the maximal tori $T_{1}$ and $T_{2}$ of two closed connected subgroups $H_{1}$ and $H_{2}$ to be contained in $T$ by passing to conjugates of $H_{1}$ and $H_{2}$ if necessary.
2.3 Enlarging tori while keeping intersections almost intact
The ranks of two closed subgroups $H_{1}$ and $H_{2}$ of a compact Lie group $G$ are clearly bounded by the rank of $G$ in the sense that
$\operatorname{rank}{G}\geq\operatorname{rank}{H_{1}}+\operatorname{rank}{H_{2}}-\mathrm{interrank}(H_{1},H_{2})$
The following proposition shows that, when $H_{1}$ and $H_{2}$ are tori, we can always enlarge $H_{2}$ such that the above inequality becomes an equality, but without affecting the intersection rank.
Proposition 29.
Consider a compact connected Lie group $G$ with a maximal torus $T$. Given two subtori $T_{1},T_{2}\subseteq T$, there exists a torus $T_{1}^{+}$ intermediate between $T_{1}$ and $T$ such that
$\operatorname{rank}G=\operatorname{rank}T_{1}^{+}+\operatorname{rank}T_{2}-\mathrm{interrank}(T_{1}^{+},T_{2})\quad\text{ and}$
$\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2}).$
Remark 30.
Even when the triple $(G,T_{1},T_{2})$ satisfies the strict biquotient condition, it is not in general possible to find a torus $T_{1}^{+}$ satisfying the above conditions and such that $(G,T_{1}^{+},T_{2})$ still satisfies the biquotient condition. In passing from $T_{1}$ to $T_{1}^{+}$, it can happen that the intersections $T_{1}^{+}\cap w(T_{2})$ acquire additional, non-central points. For example, consider the rank-two group $G:=\mathrm{SU}(3)$ with its standard diagonal torus $T=\{\operatorname{diag}(a,b,a^{-1}b^{-1})\}$, the trivial subtorus $T_{1}:=\{\operatorname{diag}(1,1,1)\}$ and the rank-one subtorus $T_{2}:=\{\operatorname{diag}(a,a^{-1},1)\}$ defined by one of the simple coroots. Clearly, the triple $(G,T_{1},T_{2})$ satisfies the strict biquotient condition. However, any rank-one subtorus $T_{1}^{+}\subseteq T$ intersects at least one of the conjugates $T_{2}$, $T_{2}^{\prime}:=\{\operatorname{diag}(1,a,a^{-1})\}$ and $T_{2}^{\prime\prime}:=\{\operatorname{diag}(a,1,a^{-1})\}$ of $T_{2}$ non-trivially, in a non-central point, so that $(G,T_{1}^{+},T_{2})$ does not even satisfy the lax biquotient condition.
Proof.
Let us abbreviate the rank as follows: $r:=\operatorname{rank}T$, $r_{1}:=\operatorname{rank}T_{1}$, $r_{2}:=\operatorname{rank}T_{2}$, $r^{\prime}:=\mathrm{interrank}(T_{1},T_{2})$. By Proposition 28, $r^{\prime}=\max_{w\in\mathrm{W}}\operatorname{rank}{(T_{1}\cap w(T_{2}))}$. In view of Lemma 27, it suffices to construct an intermediate torus $T_{1}^{+}$ of rank $r_{1}^{+}:=r-r_{2}+r^{\prime}$ such that $\operatorname{rank}{(T_{1}^{+}\cap w(T_{2}))}=r^{\prime}$ for each of the finitely many Weyl group elements $w\in\mathrm{W}$.
The proof works by translating the statement from subgroups of $T$ to quotients of $\mathbb{Z}^{r}$, to subgroups of $\mathbb{Z}^{r}$ and finally to sub-vector spaces of $\mathbb{Q}^{r}$, where it can be solved essentially by linear algebra. The following paragraphs provide some details. To begin with, we observe:
Let $W_{1}$ and $W_{2}$ be sub-vector spaces of $\mathbb{Q}^{r}$, with $c_{i}:=\dim W_{i}$. For any integer $c_{1}^{-}$ such that $0\leq c_{1}^{-}\leq\dim(W_{1}+W_{2})-c_{2}$, there exists a sub-vector space $W_{1}^{-}\subseteq W_{1}$ of dimension $c_{1}^{-}$ such that $\dim(W_{1}^{-}+W_{2})=c_{1}^{-}+c_{2}$. Moreover, the sub-vector spaces $W_{1}^{-}$ with these properties form a (non-empty) Zariski-open subset of rational Grassmannian $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$.
This observation could be phrased more elegantly, and proved, by noting that $\dim(W_{1}+W_{2})-c_{2}=c_{1}-\dim(W_{1}\cap W_{2})$, and by noting that the condition on the dimension of $W_{1}^{-}+W_{2}$ is equivalent to the condition $W_{1}^{-}\cap W_{2}=\{0\}$. Such $W_{1}^{-}$ exists because we can simply choose a complement $W^{\prime}$ of $W_{1}\cap W_{2}$ in $W_{1}$, and then pick an arbitrary subspace $W_{1}^{-}\subseteq W^{\prime}$ of dimension $c_{1}^{-}$. To see that trivial intersection with $W_{2}$ is an open condition on $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$, note that this condition on $W_{1}^{-}$ can be rephrased as: the restriction of the natural projection $W_{1}\twoheadrightarrow W_{1}/(W_{1}\cap W_{2})$ to $W_{1}^{-}$ is a monomorphism. By change of coordinates, we may assume without loss of generality that this projection is the projection onto the first $c_{1}^{-}$ coordinates. Then the restriction is a monomorphism if and only if, for any basis $(b_{1},\dots,b_{c_{1}^{-}})$ of $W_{1}^{-}$, the leading principal $(c_{1}^{-}\times c_{1}^{-})$-minor is non-zero, which is clearly an open condition on the Grassmannian.
Next, we replace the single space $W_{2}$ by finitely many sub-vector spaces $W_{2}^{(i)}$:
Suppose $W_{1}$ and $W_{2}^{(1)},\dots,W_{2}^{(n)}$ are sub-vector spaces of $\mathbb{Q}^{r}$, with all $W_{2}^{(i)}$ of equal dimension. Let $c_{1}:=\dim W_{1}$, $c_{2}:=\dim W_{2}^{(i)}$, $c^{\prime}:=\min_{i}\dim(W_{1}+W_{2}^{(i)})$ and $c_{1}^{-}:=c^{\prime}-c_{2}$. There exists a sub-vector space $W_{1}^{-}\subseteq W_{1}$ such that $c_{1}^{-}=\dim W_{1}^{-}$ and, for each $i$, $c^{\prime}=\dim(W_{1}^{-}+W_{2}^{(i)})$.
Indeed, for each $i$, we have $c_{2}\leq c^{\prime}\leq\dim(W_{1}+W_{2}^{(i)})$, so $0\leq c_{1}^{-}\leq\dim(W_{1}+W_{2}^{(i)})-c_{2}$. Thus, by the previous observation, for each $i$ the sub-vector spaces $W_{1}^{-}\subseteq W_{1}$ with $\dim(W_{1}^{-}+W_{2}^{(i)})=c^{\prime}$ form a non-empty Zariski-open subset of $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$. So we can pick any $W_{1}^{-}$ that lies in the intersection of these finitely many Zariski open subsets.
From the above observation concerning sub-vector spaces of $\mathbb{Q}^{r}$, we deduce an analogous observation concerning direct summands of $\mathbb{Z}^{r}$:
Suppose $K_{1}$ and $K_{2}^{(1)},\dots,K_{2}^{(n)}$ are direct summands of $\mathbb{Z}^{r}$, with all $K_{2}^{(i)}$ of equal rank. Let $c_{1}:=\operatorname{rank}K_{1}$, $c_{2}:=\operatorname{rank}K_{2}^{(i)}$, $c^{\prime}:=\min_{i}\operatorname{rank}(K_{1}+K_{2}^{(i)})$ and $c_{1}^{-}:=c^{\prime}-c_{2}$. There exists a direct summand $K_{1}^{-}$ of $\mathbb{Z}^{r}$ such that $K_{1}^{-}\subseteq K_{1}$ with $c_{1}^{-}=\operatorname{rank}K_{1}^{-}$ and, for each $i$, $c^{\prime}=\operatorname{rank}(K_{1}^{-}+K_{2}^{(i)})$.
Indeed, note that for a sub-vector space $W\subseteq\mathbb{Q}^{r}$, the intersection $W\cap\mathbb{Z}^{r}$ is a direct summand of $\mathbb{Z}^{r}$, of rank equal to the dimension of $W$. Conversely, from any subgroup $K\subseteq\mathbb{Z}^{r}$, we can obtain the sub-vector space $K\otimes_{\mathbb{Z}}\mathbb{Q}\subseteq\mathbb{Q}^{r}$ of dimension equal to the rank of $K$. We obtain in this way a bijection between sub-vector spaces of $\mathbb{Q}^{r}$ and direct summands of $\mathbb{Z}^{r}$. Moreover, for arbitrary subgroups $K_{1},K_{2}\subseteq\mathbb{Z}^{r}$, we have $(K_{1}+K_{2})\otimes\mathbb{Q}=K_{1}\otimes\mathbb{Q}+K_{2}\otimes\mathbb{Q}$. (Note that $K_{1}+K_{2}$ may not be a direct summand even when $K_{1}$ and $K_{2}$ are.) Thus, taking the vector spaces $W_{1}:=K_{1}\otimes\mathbb{Q}$ and $W_{2}^{(i)}:=K_{2}^{(i)}\otimes\mathbb{Q}$ in the previous observation, we obtain a vector space $W_{1}^{-}\subseteq W_{1}$ with the properties specified above, and we may take $K_{1}^{-}:=W_{1}^{-}\cap\mathbb{Z}^{r}$. As noted, this is a direct summand of $\mathbb{Z}^{r}$, clearly contained in $K_{1}$, and the ranks work out as intended. Finally, using Lemma 4, the observation corncerning direct summands of $\mathbb{Z}^{r}$ may be translated into the following statement about tori:
Suppose $T_{1}$ and $T_{2}^{(1)},\dots,T_{2}^{(n)}$ are subtori of a torus $T$, with all $T_{2}^{(i)}$ of equal rank. Let $r:=\operatorname{rank}T$, $r_{1}:=\operatorname{rank}K_{1}$, $r_{2}:=\operatorname{rank}K_{2}^{(i)}$, $r^{\prime}:=\max_{i}\operatorname{rank}(T_{1}\cap T_{2}^{(i)})$ and $r_{1}^{+}:=r-r_{2}+r^{\prime}$. There exists a torus $T_{1}^{+}$ such that $T_{1}\subseteq T_{1}^{+}\subseteq T$, $r_{1}^{+}=\operatorname{rank}T_{1}^{+}$ and, for each $i$, $r^{\prime}=\operatorname{rank}(T_{1}^{+}\cap T_{2}^{(i)})$.
This clearly proves the claim we set out to establish. ∎
2.4 Reduction to toroidal subgroups
Proposition 31 (tori vs. general case).
Consider a compact Lie group $G$ with connected subgroups $H_{1}$ and $H_{2}$, and suppose $T_{1}$ and $T_{2}$ are maximal tori of $H_{1}$ and $H_{2}$, respectively. If $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes for some degree $i$ and some prime ideal $\mathfrak{p}\subset\mathrm{R}G$, then so does $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}$.
In case $H_{1}$ and $H_{2}$ are good, Theorem 13 can be used to show that the implication of the proposition is even an equivalence, but we will not need this.
Proof.
By symmetry, it suffices to show how to pass from $T_{1}$ to $H_{1}$: we will argue that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}H_{2})=0$ implies $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})=0$. As in the proof [SCSQ, Corollary 3.10] (the corollary quoted as Corollary 18 above), this can be seen using holomorphic induction (Theorem 10): the inclusion $i\colon T_{1}\hookrightarrow H_{1}$ defines a homomorphism of $\mathrm{R}H_{1}$-modules $i_{*}$ splitting $i^{*}$, and homomorphisms of $\mathrm{R}H_{1}$-modules are, in particular, homomorphisms of $\mathrm{R}G$-modules. So the claim follows. Localization does not affect this argument. ∎
3 The proof
We now proceed to prove Theorem 3, starting from the most special case and generalizing in several steps. We repeatedly use the fact that $\operatorname{Tor}$ commutes with localization: for any prime ideal $\mathfrak{p}\subset\mathrm{R}G$, we have canonical isomorphisms
$[\operatorname{Tor}_{\mathrm{R}G}^{i}(\mathrm{R}H_{1},\mathrm{R}H_{2})]_{\mathfrak{p}}\cong\operatorname{Tor}_{(\mathrm{R}G)_{\mathfrak{p}}}^{i}((\mathrm{R}H_{1})_{\mathfrak{p}},(\mathrm{R}H_{2})_{\mathfrak{p}}).$ (1)
Recall from Proposition 22 (d) that $\mathrm{Z}G$ denotes the centre of $G$.
Proposition 32 (particular subgroups of maximal ranks).
Suppose $G$ is a good Lie group, and that $H_{1}$ and $H_{2}$ are connected subgroups whose ranks are maximal in the sense that
$\operatorname{rank}H_{1}+\operatorname{rank}H_{2}-\mathrm{interrank}(H_{1},H_{2})=\operatorname{rank}G.$ (2)
Assume that $H_{1}$ is a torus. Assume moreover that the restriction $\mathrm{R}G\to\mathrm{R}H_{2}$ is surjective and that its kernel is generated by a regular sequence $\mathbb{y}=(y_{1},\dots,y_{c_{2}})$ of length $c_{2}=\operatorname{rank}G-\operatorname{rank}H_{2}$. Then $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$.
Proof.
Choose a maximal torus $T\subset G$ containing the torus $H_{1}$, and let $i\colon T\hookrightarrow G$ denote the inclusion. Then $H_{1}$ is a direct factor of $T$ [HilgertNeeb]*Lem. 15.3.2, and the kernel of the restriction $\mathrm{R}T\to\mathrm{R}H_{1}$ is generated by a regular sequence $\mathbb{x}=(x_{1},\dots,x_{c_{1}})$ of length $c_{1}=\operatorname{rank}G-\operatorname{rank}H_{1}$. It follows that $\mathrm{R}H_{1}\cong\mathrm{R}T/(\mathbb{x})$ as an $\mathrm{R}T$-module, and that the Koszul complex defined by $\mathbb{x}$ provides a free resolution of $\mathrm{R}H_{1}$ as an $\mathrm{R}T$-module. We write $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})$ for this complex, and $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})\to\mathrm{R}H_{1}$ for the resolution.
For $H_{2}$, we have an isomorphism of $\mathrm{R}G$-modules $\mathrm{R}H_{2}\cong\mathrm{R}G/(\mathbb{y})$ by assumption, and the Koszul complex $\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ provides a free resolution of $\mathrm{R}H_{2}$ as an $\mathrm{R}G$-module.
Now observe that we have an isomorphism of $\mathrm{R}G$-modules, and an isomorphism of complexes of $\mathrm{R}G$-modules, as follows:
$\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\mathrm{R}H_{2}$ $\cong\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y})$ (3)
$\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})\otimes_{\mathrm{R}G}\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ $\cong\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y})$ (4)
The first isomorphism is clear. To verify the second isomorphism, we can decompose each Koszul complex into a tensor product of Koszul complexes of lengths two:
$\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})$ $\underset{\mathrm{R}T}{\cong}\bigotimes_{\begin{subarray}{c}\mathrm{R}T\\ i=1,\dots,c_{1}\end{subarray}}(\mathrm{R}T\xrightarrow{x_{i}}\mathrm{R}T)$
$\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ $\underset{\mathrm{R}G}{\cong}\bigotimes_{\begin{subarray}{c}\mathrm{R}G\\ j=1,\dots,c_{2}\end{subarray}}(\mathrm{R}G\xrightarrow{y_{j}}\mathrm{R}G)$
The isomorphism (4) can now be obtained by successively applying the isomorphisms
[…]
We have taken care to distinguish $\mathrm{R}T$-module isomorphisms from $\mathrm{R}G$-module isomorphisms in the notation, but note that every isomorphism of $\mathrm{R}T$-modules is a fortiori an isomorphism of $\mathrm{R}G$-modules.
Viewing the left side of (4) as a double complex, we obtain an associated spectral sequence of $\mathrm{R}G$-modules:
$E_{0}^{p,q}=\Lambda^{p}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})\otimes_{\mathrm{R}G}\Lambda^{q}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})\quad\Rightarrow\quad\mathrm{H}^{p+q}(\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y})),$
As $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x})$ is a resolution of $\mathrm{R}H_{1}$, and as each term of $\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ is a free $\mathrm{R}G$-module, the $E_{1}$-page is concentrated in a single column as follows:
$E^{p,q}_{1}=\begin{cases}\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\Lambda^{q}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})&\text{ for }p=0\\ 0&\text{ for }p\neq 0\end{cases}$
As $\Lambda^{\bullet}_{\mathrm{R}G}(\mathrm{R}G;\mathbb{y})$ is a free resolution of $\mathrm{R}H_{2}$, the $E_{2}$-page computes $\operatorname{Tor}$-groups as follows:
$E^{p,q}_{2}=\begin{cases}\operatorname{Tor}_{RG}^{q}(\mathrm{R}H_{1},\mathrm{R}H_{2})&\text{ for }p=0\\ 0&\text{ for }p\neq 0\end{cases}$
At this stage, the spectral sequence necessarily collapses, i. e. $E^{p,q}_{2}=E^{p,q}_{\infty}$. We thus obtain, for all $i$, an isomorphism of $\mathrm{R}G$-modules:
$\operatorname{Tor}_{\mathrm{R}G}^{i}(\mathrm{R}H_{1},\mathrm{R}H_{2})\cong\mathrm{H}^{i}(\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y}))$ (5)
The basic strategy now is to compute the Krull dimension of $\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y})$, to deduce that $(\mathbb{x},i^{*}\mathbb{y})$ is a regular sequence in $\mathrm{R}T$, and to conclude that the augmented complex $\Lambda^{\bullet}_{\mathrm{R}T}(\mathrm{R}T;\mathbb{x},i^{*}\mathbb{y})\to\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y})$ is exact, so that the cohomology groups in (5) vanish in positive degrees.
In order to implement this strategy, we now fix and localize at a prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ that lies in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$. In view of eq. 1, we may and will moreover assume that $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$.
By Proposition 22 (d), there is a unique ideal $\mathfrak{p}_{T}\subset\mathrm{R}T$ lying over $\mathfrak{p}\subset\mathrm{R}G$. It follows from the construction that $(\mathbb{x},i^{*}\mathbb{y})\subseteq\mathfrak{p}_{T}$: We have $(\mathbb{x})=\ker(\mathrm{R}T\to\mathrm{R}H_{1})$, so $(i^{*})^{-1}(\mathbb{x})\subseteq\ker(\mathrm{R}G\to\mathrm{R}H_{1})$. Moreover, by the equivalence ($b\Leftrightarrow c$) of Proposition 21, $\ker(\mathrm{R}G\to\mathrm{R}H_{1})\subseteq\mathfrak{p}$. So $(i^{*})^{-1}(\mathbb{x})\subseteq\mathfrak{p}$, and we deduce from Lemma 8 (c) that $(\mathbb{x})\subseteq\mathfrak{p}_{T}$. Similarly, $(\mathbb{y})=\ker(\mathrm{R}G\to\mathrm{R}H_{2})\subseteq\mathfrak{p}$, and as $\mathfrak{p}=(i^{*})^{-1}(\mathfrak{p}_{T})$, we find $(i^{*}\mathbb{y})\subseteq\mathfrak{p}_{T}$.
We can therefore view $(\mathbb{x},i^{*}\mathbb{y})$ as a sequence in the maximal ideal $\mathfrak{p}_{T}\cdot(\mathrm{R}T)_{\mathfrak{p}}$ of the local ring $(\mathrm{R}T)_{\mathfrak{p}}$. Using (3) and the first equality of Proposition 22 (b), we see that $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y}))$. By Lemma 12, we can find another prime $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}T/(\mathbb{x},i^{*}\mathbb{y}))$ with $\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$ such that the first equality in the following chain of equalities and inequalities holds:
$\dim(\mathrm{R}T)_{\mathfrak{p}}-\dim(\mathrm{R}T)_{\mathfrak{p}}/(\mathbb{x},i^{*}\mathbb{y})$ $=\dim(\mathrm{R}G)_{\mathfrak{p}}-\dim(\mathrm{R}G/\mathfrak{p}^{\prime})_{\mathfrak{p}}$
$=\dim(\mathrm{R}G)-\dim(\mathrm{R}G/\mathfrak{p}^{\prime})$
$\geq\operatorname{rank}G-\mathrm{interrank}(H_{1},H_{2})$
For the second equality, we use the fact that the dimension formula holds in $\mathrm{R}G$ (Lemma 11), and, a fortiori, in the localization $(\mathrm{R}G)_{\mathfrak{p}}$. The ideal $\mathfrak{p}^{\prime}$ is, in particular, contained in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$, and hence by Proposition 22 (b) it is contained in $\mathrm{Supp}_{\mathrm{R}G}(H_{1}\cap gH_{2}g^{-1})$ for some $g\in G$. The final inequality therefore follows from Proposition 22 (c) and Lemma 6 (c).
On the other hand, $(\mathbb{x},i^{*}\mathbb{y})$ is a sequence of length
$c_{1}+c_{2}$ $=(\operatorname{rank}G-\operatorname{rank}H_{1})+(\operatorname{rank}G-\operatorname{rank}H_{2})$
=(${=}\operatorname{rank}G-\mathrm{interrank}(H_{1},H_{2}).$ ) .
As $(\mathrm{R}T)_{\mathfrak{p}}$ is regular (Lemma 15), and hence in particular Cohen-Macaulay, we can now apply [Hartshorne]*Theorem 8.21A (c) or [Matsumura:CA]*§ 16.B, Theorem 31 to deduce that the $(\mathbb{x},i^{*}\mathbb{y})$ is a regular sequence. Hence, using [Eisenbud]*Corollary 17.5, we conclude that the Koszul complex $\Lambda^{\bullet}_{(\mathrm{R}T)_{\mathfrak{p}}}((\mathrm{R}T)_{\mathfrak{p}};\mathbb{x},i^{*}\mathbb{y})$ is exact in all positive degrees. ∎
Proposition 33 (tori of maximal ranks).
Suppose $G$ is a good Lie group, and that $T_{1}$ and $T_{2}$ are tori in $G$ such that
$\operatorname{rank}T_{1}+\operatorname{rank}T_{2}-\mathrm{interrank}(T_{1},T_{2})=\operatorname{rank}G.$ (6)
Then $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$.
Proof.
Consider the triple $(G\times G,T_{1}\times T_{2},\Delta G)$. The group $G\times G$ is good as $G$ is good and $\mathrm{R}(G\times G)\cong\mathrm{R}(G)\otimes\mathrm{R}(G)$. The subgroups $T_{1}\times T_{2}$ and $\Delta G$ are connected. Using Proposition 25, we see that the triple still satisfies the maximal rank condition (2). The restriction $\Delta^{*}\colon\mathrm{R}(G\times G)\to\mathrm{R}(G)$ is surjective. Also, given that $G$ is good, we can easily find a regular sequence of length $\operatorname{rank}G$ generating the kernel of $\Delta^{*}$. So Proposition 32 applies to $(G\times G,T_{1}\times T_{2},\Delta G)$. Using Proposition 25, we conclude that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$. ∎
Proposition 34 (arbitrary subgroups).
Suppose $G$ is a good Lie group, and that $H_{1}$ and $H_{2}$ are connected subgroups. Then $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$ for all
$i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})+\mathrm{interrank}(H_{1},H_{2})$
and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$.
Proof.
Choose maximal tori $T_{1}$ and $T_{2}$ of $H_{1}$ and $H_{2}$, respectively. Replacing $H_{1}$ and $H_{2}$ by conjugate subgroups if necessary, we may assume that $T_{1}$ and $T_{2}$ are contained in a joint maximal torus $T$ of $G$. Note that by Proposition 26, passing from the groups $H_{i}$ to their maximal tori $T_{i}$ does not change the intersection rank. By Proposition 29, we can find a torus $T_{1}^{+}$ such that $T_{1}\subseteq T_{1}^{+}\subseteq T$, such that $\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2})$ and such that the triple $(G,T_{1}^{+},T_{2})$ satisfies the maximal rank condition (6). So by Proposition 33, $\operatorname{Tor}_{*}^{\mathrm{R}G}(\mathrm{R}(T_{1}^{+}),\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes in all positive degrees for all $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$
We can now apply Theorem 19 with $\Lambda$, $\Gamma$, $A$ and $C$ as follows:
${(\mathrm{R}T_{2})_{\mathfrak{p}}}$${(\mathrm{R}T_{1})_{\mathfrak{p}}}$${(\mathrm{R}G)_{\mathfrak{p}}}$${(\mathrm{R}T_{1}^{+})_{\mathfrak{p}}}$
This gives us isomorphisms
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{2},\mathrm{R}T_{1})_{\mathfrak{p}}\cong\operatorname{Tor}_{i}^{\mathrm{R}T_{1}^{+}}(\dots,\mathrm{R}T_{1})_{\mathfrak{p}}$
in all degrees $i$. The module on the right vanishes for all $i>\operatorname{pd}_{\mathrm{R}T_{1}^{+}}\mathrm{R}T_{1}$, and by Corollary 18, $\operatorname{pd}_{\mathrm{R}T_{1}^{+}}\mathrm{R}T_{1}\leq\operatorname{rank}T_{1}^{+}-\operatorname{rank}T_{1}$. As $(G,T_{1}^{+},T_{2})$ satisfies (6), and as $\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2})$, this bound can be rewritten as $\operatorname{rank}T_{1}^{+}-\operatorname{rank}T_{1}=\operatorname{rank}G-\operatorname{rank}T_{1}-\operatorname{rank}T_{2}+\mathrm{interrank}(T_{1},T_{2})$. We thus deduce that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes in the required range. Using Proposition 31, we obtain the same claim for the triple $(G,H_{1},H_{2})$. ∎
Theorem 3 is now just a special case:
Proof of Theorem 3.
Fix some degree $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})+\mathrm{interrank}(H_{1},H_{2})$. Any prime $\mathfrak{p}$ supporting $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ is necessarily contained in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$. So by Proposition 22 (b), it is contained in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(H_{1}\cap gH_{2}g^{-1})$ for some $g\in G$. As $(G,H_{1},H_{2})$ satisfies the lax biquotient condition, each of the intersections $H_{1}\cap gH_{2}g^{-1}$ is contained in the centre $\mathrm{Z}G$. So by Proposition 22 (a) we have $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$, hence by Proposition 34, $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$. ∎
Finally, Theorem 1 is, of course, a special case of Theorem 3.
4 K-theory of biquotients
This short final section explains how to apply Theorem 1 to biquotients to obtain Corollary 2.
[…]
Suppose $G$ is a compact connected Lie group with torsion-free fundamental group. Then for $G$-spaces $X$ and $Y$ we have a strongly convergent spectral sequence
$\{\operatorname{Tor}^{p}_{\mathrm{R}G}(\mathrm{K}_{G}^{*}(X),\mathrm{K}_{G}^{*}(Y))\}^{q}\Rightarrow\mathrm{K}^{p+q}_{G}(X\times Y).$
Here, the $\mathrm{R}G$-modules $\mathrm{K}_{G}^{*}(X)$ and $\mathrm{K}_{G}^{*}(Y)$ are understood to be $\mathbb{Z}/2$-graded, and $\operatorname{Tor}$ denotes graded Tor. The $G$-spaces are assumed to be compactly generated and have the homotopy type of a CW complex, see [Hodgkin, § I.1].
Given a strict biquotient $H_{1}\backslash G/H_{2}$, we have a principal $G$-bundle $H_{1}\backslash G\times G/H_{2}\to H_{1}\backslash G/H_{2}$ [Singhof:DCM]. Take $X:=H_{1}\backslash G$ and $Y:=G/H_{2}$. Then $\mathrm{K}^{*}_{G}(X)\cong\mathrm{R}H_{1}$, $\mathrm{K}^{*}_{G}(Y)\cong\mathrm{R}H_{2}$, and $\mathrm{K}^{*}_{G}(X\times Y)\cong\mathrm{K}^{*}(H_{1}\backslash G/H_{2})$. So the spectral sequence takes the form
$\operatorname{Tor}^{p}_{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})\Rightarrow\mathrm{K}^{p}(H_{1}\backslash G/H_{2}).$
In particular, given that $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ are concentrated in degree zero, the graded Tor is just the usual Tor. Theorem 1 therefore implies the results on the K-theory of biquotients summarized in Corollary 2 in exactly the same way that Steinberg’s Theorem and more generally Corollary 18 provide a computation of the K-theory of homogeneous spaces. We refer to [SCSQ, proof of Theorem 3.6 and, in particular, Figure 1] for details of the argument.